OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 7x - 14 = 6x + 2 Equation #2 7 - 2x = 5x + 21 A) The solution to equation #1 is smaller than the solution to equation #2. B) The solution to equation #1 is larger than the solution to equation #2. C) The solution to equation #1 is the same as the solution to equation #2. D) None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

1) 7x – 14 = 6x + 2 7x-6x= 2+14 x= 16 7 – 2x = 5x + 21 7-21= 5x+2x -14=7x x= -2 So the answer is B 2) 6x + 12 = 4(x + 6) 6x+12= 4x+24 6x-4x= 24-12 2x= 12 x= 6

OpenStudy (anonymous):

So its B?

OpenStudy (anonymous):

so the answer is A

OpenStudy (anonymous):

7x - 14 = 6x + 2 7x - 6x = 2 + 14 x = 16 7 - 2x = 5x + 21 7 - 21 = 5x + 2x - 14 = 7x -2 = x The answer is B.

OpenStudy (anonymous):

yeah the other one is.

OpenStudy (anonymous):

what other one ?

OpenStudy (anonymous):

nevermind-________-